# Latent Diffusion Model-Enabled Real-Time Semantic Communication Considering Semantic Ambiguities and Channel Noises

### v1

@@ Abstract
Semantic communication (SemCom) has emerged as a new paradigm for communication systems, with deep learning (DL) models being one of the key drives to shift from the accuracy of bit/symbol to the semantics and pragmatics of data. Nevertheless, DL-based SemCom systems often face performance bottlenecks due to overfitting, poor generalization, and sensitivity to outliers. Furthermore, the varying-fading gains and noises with uncertain signal-to-noise ratios (SNRs) commonly present in wireless channels usually restrict the accuracy of semantic information transmission. Consequently, to address the aforementioned issues, this paper constructs a SemCom system based on the latent diffusion model, and proposes three improvements compared to existing works: i) To handle potential outliers in the source data, semantic errors obtained by projected gradient descent based on the vulnerabilities of DL models, are utilized to update the parameters and obtain an outlier-robust encoder. ii) A lightweight single-layer latent space transformation adapter completes one-shot learning at transmitter and is placed before the decoder at receiver, enabling adaptation for out-of-distribution data or enhancing human-perceptual quality. iii) An end-to-end consistency distillation (EECD) strategy is used to distill the diffusion models trained in latent space, enabling deterministic single or few-step real-time denoising in various noisy channels while maintaining high semantic quality. Extensive numerical experiments across different datasets demonstrate the superiority of the proposed SemCom system, consistently proving its robustness to outliers, the capability to transmit data with unknown distributions, and the ability to perform real-time channel denoising tasks while preserving high human perceptual quality, outperforming the existing denoising approaches in semantic metrics such as MS-SSIM and LPIPS.
@@ I Introduction
With the booming development of artifical intelligence (AI), augmented and virtual reality [1], 4K/6K straming [2], and the intelligent sensing devices for smart grids [3] and vehicles [4] within the internet of things (IoT), an efficient and reliable communication system becomes an essential component in the realm of 6-th generation (6G) communications [5]. In information and communication technology, joint source-channel coding (JSCC) [6, 7] is committed to the integrated design of source and channel codes for efficient transmission of data, leveraging Shannon information theory. However, classic JSCC techniques, employing coding methods for engineering applications such as WebP [8], JPEG [9], JPEG2000 [10], and BPG [11], have solely focused on the statistical characteristics of the data being transmitted, disregarding the semantic content they encompass.
@@
Recently, the pursuit of more efficient and intelligent data transmission has given rise to semantic communication (SemCom) systems [12], where the focus is shifted from traditional bit-level accuracy to the conveyance of meaning and intent. The essence of SemCom lies in its capacity to prioritize the transmission of essential information, thus promising significant improvements in bandwidth utilization and overall communication efficiency [13]. Fortunately, with the rapid advancement of machine learning, deep learning (DL) based SemComs have emerged as an improtant approach for the implementation of SemCom [14]. Specifically, SemCom built upon neural networks such as variational autoencoder (VAE) [15], residual network (ResNet) [16], convolutional neural network (CNN) [17], long short-term memory (LSTM) network [18], generative adversarial network (GAN) [19], and Transformer [14] have demonstrated effectiveness in extracting the semantic features of source data. This allows for the mapping of source data into a lower-dimensional space for transmission over noisy wireless channels to the receiver, where it can ultimately be decoded back into its original form, whether that be images [20], audio [21], text [14], or multimodal data [22]. Nonetheless, the intrinsic complexity of semantic information, coupled with the unpredictable nature of communication channels, poses new challenges that those SemCom systems are not designed to handle.
@@
Currently, diffusion models (DMs) have taken the forefront in the field of AI-generated content (AIGC) and have achieved remarkable and landmark advancements [23, 24], surpassing other generative models such as VAE [25], GAN [26], and Normalizing Flow (NF) [27]. Consequently, the application of DMs to tackle challenges within SemCom systems is beginning to gain traction [28]. Conditional DM, guided by semantic and free-space information from other users, progressively generate matching data for mixed visual reality applications [29]. Similarly, conditional DMs, guided by invertible neural networks [30], compressed one-hot maps [31], decoded low-quality data [32], and scene graphs [33], have been proposed for image transmission to achieve higher perceptual quality. DMs have also been adapted to rectify errors caused by varying-fading gains and low signal-to-noise ratio (SNR) noises [34]. Wireless channel estimation has also been performed by complex architectures based on DMs [35, 36], whose practicality issues due to high complexity remain to be addressed. Besides serving as decoders for JSCC, DMs can also act as denoisers placed after decoders to enhance data quality [37]. In [38] and [39], prompts, latent embeddings, or noisy data are transmitted over wireless channels to the receiver as starting points or input conditions for DMs, inevitably increasing bandwidth burden. However, the primary bottleneck of DMs lies in their slow data generation speed due to the multi-step prediction process required to improve generation quality, making such time-consuming communication impractical for real-time SemCom and edge users. Thus, some denoising or encoding methods opt for latent DM (LDM) [40] or acceleration techniques [41, 42] to significantly reduce the computational complexity. Nevertheless, since these enhanced approaches still feature a multi-step process during sampling, they inadequately address the challenges of real-time SemCom.
@@
DMs-based SemCom offers high perceptual quality but also introduces a bottleneck of high latency. Moreover, structural errors or noises and data following unknown distributions can introduce inaccuracies and distortions in the transmitted information when DL-based SemCom systems are deployed. The former, known as semantic errors [43], can arise from exploiting the vulnerabilities of DL models by adversarial attacks that lead to semantic discrepancies. Additionally, when DL-based SemCom system trained by the specific category of data transmits out-of-distribution data [44, 19], the reconstructed data at the receiver may also be semantically ambiguous due to the prevalent issues of poor generalization and overfitting in current AI models. To balance generation quality with speed, LDMs have been chosen as the underlying proposed architecture for the SemCom approach. In summary, while LDMs excel at abstracting and encoding semantic content, the transmission process remains susceptible to semantic ambiguities from semantic errors or out-of-distribution data, i.e., misinterpretations of the meaning or intention behind the transmitted data, and channel noises under different conditions.
@@
To address these issues, this paper presents a comprehensive framework that enhances real-time SemCom by leveraging the capabilities of LDMs while simultaneously considering the effects of semantic ambiguities and channel imperfections. The proposed SemCom model builds upon and enhances the foundational architecture of a pretrained Wasserstein GAN [45] with VAE (VAE-WGAN). The overall contribution of this approach is threefold:
@@
    Semantic errors caused by measurement inaccuracies and cyber attacks can significantly disrupt the normal encoding and decoding of deep JSCC systems. To address this, the vulnerabilities of the pretrained encoder and generator are exploited using convex optimization to determine the most significant undetectable semantic errors. The pretrained encoder is then updated with the obtained semantic errors to refine the neural network parameters, making the encoder robust and resilient to anomalously transmitted data. This parameter update process with data augmentation is self-supervised.
@@
    A rapid domain adaptation strategy is introduced to ensure the reconstructed data is semantically accurate at the receiver when SemCom system transmits data with an unknown distribution. This strategy employs two additional lightweight single-layer neural networks that perform online one-shot or few-shot learning based on adversarial learning strategies. The updated parameters are transmitted to the dynamic neural network deployed at the receiver through the shared knowledge of SemCom system, while the parameters of other networks remain unchanged, thus achieving low-cost out-of-distribution latent space transformation.
@@
    Inspired by channel denoising DMs [40] and consistency distillation [46], the LDM based on ordinary differential equation (OED) trajectories and variance explosion strategy is trained with known channel state information (CSI) [35, 36]. During the sampling phase, it can denoise the received equalized signals according to different CSIs. Furthermore, the end-to-end consistency distillation (EECD) approach that considers semantic metrics is proposed to distill the trained LDM, ultimately transforming the multi-step denoising process into a deterministic one-step real-time denoising procedure, capable of flexibly addressing varying fading channels and uncertain SNRs.
@@
The efficiency and reliability of the proposed SemCom system in term of perceptual quality and timeliness are validated by rigorous and extensive experiments, providing concrete evidence of its superiority over conventional methods.
@@
The rest of this paper is organized as follows. Section II briefly introduces the proposed wireless SemCom system model and existing challenges. Section III elaborates on the JSCC design of the proposed SemCom system for unknown errors and distributions. The real-time channel denoising implementation is established by EECD in Section IV. Numerical experiments are given in Section V. Section VI concludes the paper. Supporting lemmas are included in the Appendix for reference.
@@
Figure 1: The proposed semantic communication system model with three addressed deep learning-based communication challenges: ① robust GAN inversion with semantic errors, ② domain adaptation with unknown distribution, and ③ real-time wireless channel denoising with end-to-end consistency distillation.
II System Model and Problem Formulation
@@
The proposed SemCom system model depicted in Fig. 1 consists of the encoder $E_{\bm{\phi}}(.)$ with target distribution $q_{\bm{\phi}}(\bm{z}|\bm{x})$, diffusion model $\bm{\epsilon}_{\bm{\theta}}(.,.)$ with denoised latent vector’s distribution $p_{\bm{\theta}}(\bm{z})$, and decoder $G_{\bm{\psi}}(.)$ with reconstruction target distribution $q_{\bm{\psi}}(\bm{x}|\bm{z})$ by utilizing the synthesized encodings $\bm{z}$ from diffusion model. The goal of training this LDM is to learn $\left\{\bm{\phi},\bm{\theta},\bm{\psi}\right\}$ by minimizing the overall variational upper bound [47]
@@
$=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[-\log q_{\bm{\psi}}(\bm{x}|\bm{z})\right]+\mathbb{E}_{q}\left[\mathcal{D}_{KL}\left(q_{\bm{\phi}}(\bm{z}|\bm{x})\parallel p_{\bm{\theta}}(\bm{z})\right)\right]$
@@
where $q_{\bm{\phi}}(\bm{z}|\bm{x})$ approximates the true posterior $q_{\bm{\psi}}(\bm{z}|\bm{x})$. The loss in Eq. (1) has been widely applied and validated in fast data generation [48]. Unlike data generation, the goal of SemCom systems is to make the reconstructed data in receivers shows desired meaning. Consequently, Eq. (1) is divided into three terms: the encoding entropy term at the transmitter, the corss entropy term at the wireless channel, and the reconstruction term at the receiver. However, as stated in Section I, the LDM-enabled SemCom system still faces the following three major challenges:
1) Semantic Error: Due to unreasonable measurement, shooting, storage methods, or cyber attacks, the transmitted data may contain some imperceptible errors or noises, which may cause DL-based communication systems to reconstruct data with wrong semantics based on the contaminated data $\bm{x}^{\prime}$ at the receiver. 2) Unknown Distribution: When a DL-based communication system transmits data $\bm{x}^{\prime\prime}$ with unknown distribution, i.e., the data type is not included in the training dataset, the decoder may generate data with different meaning. A lot of works on transfer learning are dedicated to address the challenge of domain adaptation [19, 44, 49], but strategies for quickly transmit out-of-distribution data still need further exploration. 3) Channel Uncertainties: The wireless channels are inevitably subjected to varying fading gains and noises with uncertain SNRs. For this reason, the received latent bottleneck is the data $\bm{z}^{\prime}$ with noises and bit errors. Diffusion models can utilize the cross entropy term $\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[-\log p_{\bm{\theta}}(\bm{z})\right]$ to solve the mapping $p(\bm{z}|\bm{z}^{\prime})$, correct bit errors [34], or do joint channel estiamtion and denoising [35, 36, 40]. Nontheless, these approaches still require uncertain multi-step evaluations to generate clean $\bm{z}$, hindering the real-time implementation of SemCom systems. As a result, the communication objective terms in Eq. (1) are rewritten as:
@@
The proposed SemCom system solves the potential problems of DL-based communication system point-to-point. The basic encoder-decoder architecture of the proposed system consists of a variational encoder $E_{\bm{\phi}}(.)$ and generator $G_{\bm{\psi}}(.)$ of WGAN. The fundamental principles and joint training algorithm are detailed in Subsection III-A and Subsection III-B. As illustrated in Fig. 1, the threefold improvements are further clarified as follows:
1) Robust GAN Inversion: The imperceptible semantic error that leads to the maximum reconstruction error in DL-based SemCom systems is defined and obtained through adversarial convex optimization. Based on this semantic error, the parameters of the optimized robust encoder are updated from $\bm{\phi}$ to $\bm{\phi}^{\prime}$ to encode normal latent space for transmission. The specific GAN robust inversion method is detailed in Subsection III-C. 2) Domain Adaptation: When transmitting out-of-distribution data, the lightweight single-layer $g_{\bm{\omega}}$ and $d_{\bm{\nu}}$ are exploited for one-shot fast and adversarial domain adaptation learning. The learned parameter $\bm{\omega}$ will be simultaneously transmitted to the receiver along with the data for latent space transformation, and the decoder will ultimately output semantically consistent data. The specific implementation can be found in Subsection III-D. 3) Real-time Channel Denoising: Assume that the CSIs are known, EECD is proposed to distill LDM from multi-step denoising process into one-setp, thereby reducing the computational complexity of online sampling during real-time communication. The detailed wireless channel modeling, training and sampling approaches of the latent channel denoising DM, and one-step real-time channel denoising algorithm are elucidated in Section IV.
III Deep JSCC for unknown errors and distributions

### v4

# Latent Diffusion Model-Enabled Low-Latency Semantic Communication in the Presence of Semantic Ambiguities and Wireless Channel Noises

@@ Abstract
Deep learning (DL)-based Semantic Communications (SemCom) is becoming critical to maximize overall efficiency of communication networks. Nevertheless, SemCom is sensitive to wireless channel uncertainties, source outliers, and suffer from poor generalization bottlenecks. To address the mentioned challenges, this paper develops a latent diffusion model-enabled SemCom system with three key contributions, i.e., i) to handle potential outliers in the source data, semantic errors obtained by projected gradient descent based on the vulnerabilities of DL models, are utilized to update the parameters and obtain an outlier-robust encoder, ii) a lightweight single-layer latent space transformation adapter completes one-shot learning at the transmitter and is placed before the decoder at the receiver, enabling adaptation for out-of-distribution data and enhancing human-perceptual quality, and iii) an end-to-end consistency distillation (EECD) strategy is used to distill the diffusion models trained in latent space, enabling deterministic single or few-step low-latency denoising in various noisy channels while maintaining high semantic quality. Extensive numerical experiments across different datasets demonstrate the superiority of the proposed SemCom system, consistently proving its robustness to outliers, the capability to transmit data with unknown distributions, and the ability to perform real-time channel denoising tasks while preserving high human perceptual quality, outperforming the existing denoising approaches in semantic metrics such as multi-scale structural similarity index measure (MS-SSIM) and learned perceptual image path similarity (LPIPS).
@@ I Introduction
With the booming development of artifical intelligence (AI), augmented and virtual reality [1], 4K/6K straming [2], and the intelligent sensing devices for smart grids [3] and vehicles [4] within the internet of things (IoT), an efficient and reliable communication system becomes an essential component in the realm of 6-th generation (6G) communications [5]. In information and communication technology, joint source-channel coding (JSCC) [6] is committed to the integrated design of source and channel codes for efficient transmission of data, leveraging Shannon information theory. However, classic JSCC techniques, employing coding methods for engineering applications such as JPEG [7], JPEG2000 [8], and BPG [9], have solely focused on the statistical characteristics of the data being transmitted, disregarding the semantic content they encompass.
@@
Recently, the pursuit of more efficient and intelligent feature extraction and data transmission has given rise to semantic communication (SemCom) systems [10], where the focus has shifted from traditional bit-level accuracy to the conveyance of meaning and intent. The essence of SemCom lies in its capacity to emphasize the transmission of semantic information, thus promising significant improvements in bandwidth utilization and overall communication efficiency [11]. Fortunately, with the rapid advancement of machine learning, deep learning (DL)-based SemCom systems are becoming crucial [12]. Specifically, SemCom built upon neural networks such as variational autoencoder (VAE) [13], residual network (ResNet) [14], convolutional neural network (CNN) [15], long short-term memory (LSTM) network [16], generative adversarial network (GAN) [17], and Transformer [12] have demonstrated effectiveness in extracting the semantic features of source data. This allows for the mapping of source data into a lower-dimensional space for transmission over noisy wireless channels to the receiver, where it can ultimately be decoded back into its original form, whether that be images [18], audio [19], text [12], or multimodal data [20]. Nonetheless, the intrinsic complexity of semantic information, coupled with the uncertainity of communication channels, poses new challenges that some SemCom systems are not designed to handle.
@@
Diffusion models (DMs) have taken the forefront in the field of AI-generated content (AIGC) and have achieved remarkable advancements in generation quality [21, 22], surpassing other generative models in recent years. Consequently, the application of DMs to tackle challenges within SemCom systems is beginning to gain attraction [23, 24]. Conditional DM, guided by semantic information from other users, progressively generate matching data for mixed reality applications [25]. Similarly, conditional DMs, guided by invertible neural networks [26], compressed one-hot maps [27], decoded low-quality data [28], and scene graphs [29], have been proposed for image transmission to achieve higher perceptual quality. DMs have also been adapted to rectify errors caused by channels with varying-fading gains and low signal-to-noise ratio (SNR) noises [30]. Moreover, wireless channel estimation has also been performed by well-designed complex architectures based on DMs [31, 32]. Besides serving as decoders for joint source-channel coding (JSCC) [6], DMs can also act as denoisers placed after decoders to enhance data quality [33]. In [34] and [35], prompts, latent embeddings, or noisy data are transmitted over wireless channels to the receiver as starting points or input conditions for high-quality reverse process of DMs, inevitably increasing bandwidth burden. However, the primary bottleneck of DMs lies in their slow data generation speed due to the multi-step process in original high-dimensional pixel space required to improve reconstruction quality, making such time-consuming communication impractical for ultra low-latency SemCom and edge users. Thus, some denoising or encoding methods opt for latent DM (LDM) [36] or acceleration techniques [37, 38] to significantly reduce the computational complexity by only sampling in low-dimensional latent space. Nevertheless, since these enhanced approaches [36, 37, 38] still feature a multi-step process during sampling, they inadequately address the challenges of real-time SemCom.
@@
DMs and LDMs-based SemCom systems offer high perceptual quality but also introduces a bottleneck of high-latency. Moreover, structural errors, noises, and data following unknown distributions can introduce inaccuracies and distortions in the transmitted information when DL-based SemCom systems are deployed. The former, known as semantic errors, can arise from exploiting the vulnerabilities of DL models by adversarial attacks that lead to semantic discrepancies. Additionally, when a DL-based SemCom system trained by the specific category of data transmits out-of-distribution data [17, 39], the reconstructed data at the receiver may also be semantically ambiguous due to the prevalent issues of poor generalization and overfitting in current DL models. To balance sampling quality and speed, LDM has been chosen as the underlying architecture for the SemCom approach due to its excellent semantic encoding, semantic decoding, and channel denoising capabilities [40]. Overall, the LDM-enabled SemCom also remains susceptible to semantic ambiguities from outliers or out-of-distribution data. Furthermore, when faced with noisy wireless channels, the LDM-enabled channel denoising method may not meet the low-latency SemCom requirements [36].
@@
To address these issues, this paper presents a comprehensive framework that enhances low-latency SemCom by leveraging the capabilities of LDMs while simultaneously considering the effects of semantic ambiguities and channel imperfections. The proposed SemCom model builds upon and enhances the foundational architecture of a pretrained Wasserstein GAN [41] with VAE (VAE-WGAN). The overall contribution of this approach is threefold:
@@
    Semantic errors can significantly disrupt the normal semantic encoding and decoding of DL-based JSCC systems. To address this, the vulnerabilities of the pretrained encoder and generator are exploited using convex optimization to determine the most significant undetectable semantic errors. The pretrained encoder is then updated with the obtained semantic errors to refine the neural network parameters, making the encoder robust and resilient to anomalously transmitted data. This parameter update process with data augmentation is self-supervised.
@@
    A rapid domain adaptation strategy is introduced to ensure the reconstructed data is semantically accurate at the receiver when the SemCom system transmits data with an unknown distribution. This strategy employs two additional lightweight single-layer neural networks that perform online one-shot or few-shot learning based on adversarial learning strategies. The updated parameters are transmitted to the dynamic neural network deployed at the receiver through the shared knowledge [10] of the SemCom system, while the parameters of other networks remain unchanged, thus achieving low-cost latent space transformation.
@@
    Inspired by channel denoising DM (CDDM) [36] and consistency distillation [42], the LDM based on ordinary differential equation (ODE) trajectories and variance explosion strategy is trained with known channel state information (CSI) [31, 32]. During the sampling phase, it can denoise the received equalized signals according to different CSIs. Furthermore, the end-to-end consistency distillation (EECD) approach that considers semantic metrics is proposed to distill the trained LDM, ultimately transforming the multi-step denoising process into a deterministic one-step real-time denoising procedure, capable of flexibly addressing varying fading channels and uncertain SNRs.
@@
The efficiency and reliability of the proposed SemCom system in term of perceptual quality and timeliness are validated by rigorous and extensive experiments, providing concrete evidence of its superiority over conventional methods such as JPEG2000 [8] with low-density parity check (LDPC) [43], CNN-based deep JSCC [6], and CDDM [36]. The code is open-sourced at https://github.com/JianhuaPei/LDM-enabled-SemCom-system.
@@
The rest of this paper is organized as follows. Section II briefly introduces the proposed wireless SemCom system model, existing challenges, and related works. Section III elaborates on the JSCC design of the proposed SemCom system for transmitting data with unknown errors and distributions. The real-time channel denoising implementation is established by EECD in Section IV. Numerical experiments are given in Section V. Section VI concludes the paper. Supporting lemmas are included in the Appendix for reference.
II System Overview and Methodological Innovations
II-A Problem Formulation
Conventional DL-based JSCC typically consists of a semantic encoder $E_{\bm{\phi}}(\cdot)$ parameterized by $\bm{\phi}$ at the transmitter and a semantic decoder $G_{\bm{\psi}}(\cdot)$ parameterized by $\bm{\psi}$ at the receiver. The semantic encoder usually encodes the source data $\bm{x}$ into low-dimensional latent vectors $\bm{z}$ and transmits them over the wireless channel, and finally, the semantic decoder reconstructs the data based on the received signals. However, some DL-based JSCC systems face the following challenges:
  1. 1.
    Semantic Error: Due to unreasonable photographing, storage, or cyber attacks, the transmitted data may contain some imperceptible errors or noises $\bm{\delta}$, which may cause DL-based communication systems to reconstruct data with semantic ambiguities based on the contaminated data $\bm{x}^{\prime}=\bm{x}+\bm{\delta}$ at the receiver.
  2. 2.
    Unknown Distribution: When a DL-based communication system transmits data $\bm{x}^{\prime\prime}$ with unknown distribution, i.e., the data type is not included in the training dataset, the decoder may generate data with different semantics.
  3. 3.
    Channel Uncertainties: The wireless channels are inevitably subject to varying fading gains and noises with uncertain SNRs. Assume that the transmitted complex latent signal is denoted by $\bm{z}_{c}\in\mathbb{C}^{k}$, and the latent vector needs to utilize the wireless channel by $k$ times to reach the receiver, where $k$ represents the size of latent space. At time $t$, the $i$-th symbol of complex $k$-length received noisy signal $\bm{z}^{\prime}=\bm{y}_{c}$ can be represented as
    $y_{c,i}=h_{c,i}z_{c,i}+n_{c,i},$ (1)
    where $z_{c,i}$ represents the $i$-th component of $\bm{z}_{c}$, $h_{c,i}=\sum_{p=1}^{P}\alpha_{p}e^{-j2\pi f\tau_{p}(t)}$, $\alpha_{p}$ is the signal amplitude of the $p$-th path, $P$ denotes the number of paths, $f$ is the carrier frequency, $\tau_{p}(t)$ denotes the phase shift, $n_{c,i}\sim\mathcal{CN}(0,\sigma^{2})$ represents the complex Gaussian noise. Considering the effects of multipath fading and scattering, $h_{c,i}$ are independent and identically distributed (i.i.d.) Rician fading gains, which is denoted by
    $h_{c,i}=\sqrt{\frac{K}{K+1}}+\sqrt{\frac{1}{K+1}}h_{Rayleigh,i},$ (2)
    where $h_{Rayleigh,i}$ are i.i.d. Rayleigh fading gains and $K$ is the ratio of direct radio waves’ power and non-direct radio waves’ power. When $K=\infty$, the wireless channel becomes additive white Gaussian noise (AWGN) channel, and the channel becomes Rayleigh channel when $K=0$.
II-B Related Works
Existing SemCom models primarily focus on the extraction and transmission of semantic information, with few methods addressing the vulnerability of DL-based SemCom systems to semantic errors [44]. Current mainstream approaches in the fields of communication and AI for handling outliers in transmitted data still rely on anomaly detection [45] and data recovery [38]. Therefore, there is a need for training strategies for a semantic encoder that is robust to semantic errors.
The handling of out-of-distribution data in DL-based SemCom systems has been a research hotspot. Common approaches include transfer learning [17, 46], ensemble learning [47], and multi-task training [39], all of which can enhance the semantic accuracy of decoded data following unknown distribution. However, these methods face bottlenecks such as high resource demands and long processing times, so strategies for quickly transmitting out-of-distribution data still need further exploration.
In [10], the main tasks of semantic communication systems include the extraction and transmission of semantic information. Some methods focusing on extracting semantic information do not account for channel uncertainties, while those that consider channel imperfections mainly focus on the JSCC approach [6] or deploying denoisers at the receiver, such as denoising autoencoders [48], conditional GANs [49], and diffusion models [31, 36]. However, these channel uncertainty mitigation mechanisms still face issues such as high latency and low precision.
@@ II-B Related Works
Figure 1: The proposed SemCom system with three addressed DL-based communication challenges: ① robust GAN inversion with semantic errors, ② domain adaptation with unknown distribution, and ③ real-time wireless channel denoising with EECD, where $\bm{\mu}$ and $\bm{\sigma}$ are the two components of latent bottleneck of VAE, $\bm{H}_{z}$, $\bm{H}_{n}$, $\sigma^{2}$, $\bm{z}_{R}$, and $\bm{y}_{R}$ are the CSIs, real-valued transmitted encodings, and equalized received signals, respectively, as defined in Section IV. $f_{e}(\cdot)$ represents the modulation encoding for 256-QAM, while $f_{d}(\cdot)$ represents the demodulation decoding for 256-QAM. Other symbols’ definition can be found in Section II.
II-C SemCom System Overview
@@
The proposed LDM-enabled SemCom system is a JSCC approach that utilizes an additional diffusion model for channel denoising with quadrature amplitude modulation (QAM), as depicted in Fig. 1. Specifically, JSCC consists of the encoder $E_{\bm{\phi}}(\cdot)$ with target distribution $q_{\bm{\phi}}(\bm{z}|\bm{x})$ at the transmitter, DM $\bm{\epsilon}_{\bm{\theta}}(\cdot,\cdot)$ parameterized by $\bm{\theta}$ with denoised latent vector’s distribution $p_{\bm{\theta}}(\bm{z})$ at the receiver, and decoder $G_{\bm{\psi}}(\cdot)$ with reconstruction target distribution $q_{\bm{\psi}}(\bm{x}|\bm{z})$ by utilizing the synthesized encodings $\bm{z}$ from DM. The goal of training this LDM is to learn $\left\{\bm{\phi},\bm{\theta},\bm{\psi}\right\}$ by minimizing the overall variational upper bound (VUB) [50] to eliminate the gap between encodings of $E_{\bm{\phi}}(\cdot)$ and output of $\bm{\epsilon}_{\bm{\theta}}(\cdot,\cdot)$, and to ensure the quality of the decoded data $\hat{\bm{x}}$, defined as follows:
@@
$=\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[\mathcal{D}_{KL}\left(q_{\bm{\phi}}(\bm{z}|\bm{x})||p_{\bm{\theta}}(\bm{z})\right)\right]+\mathbb{E}_{q_{\bm{\phi}}(\bm{z}|\bm{x})}\left[-\log q_{\bm{\psi}}(\bm{x}|\bm{z})\right]$
@@
where $\mathcal{D}_{KL}(\cdot||\cdot)$ denotes the Kullback-Leibler divergence, and $q_{\bm{\phi}}(\bm{z}|\bm{x})$ approximates the true posterior $q_{\bm{\psi}}(\bm{z}|\bm{x})$ of decoder. The loss in Eq. (3) has been widely applied and validated in fast data generation [40]. Unlike data generation, the goal of SemCom systems is to make the reconstructed data in receivers show desired meaning. Consequently, Eq. (3) is divided into three terms: the encoding entropy term for semantic encoding at the transmitter, the cross entropy term for synthesized denoised bottlenecks $\bm{z}$ at the wireless channel, and the reconstruction term for perceptual quality at the receiver. As a result, define $\bm{x}^{\prime}/\bm{x}^{\prime\prime}$ and $\bm{z}^{\prime}$ as the transmitted data with aforementioned potential issues, the communication objective terms in Eq. (3) are rewritten as:
@@
The proposed SemCom system addresses the above three challenges of the DL-based communication system point-to-point. As detailed in Subsection III-A, the basic encoder-decoder architecture of the proposed system consists of a variational encoder $E_{\bm{\phi}}(\cdot)$ and generator $G_{\bm{\psi}}(\cdot)$ of WGAN. Based on this, as illustrated in Fig. 1, the threefold improvements are further clarified as follows:
  1. 1.
    Robust GAN Inversion: The imperceptible semantic error that leads to the maximum reconstruction error in the DL-based SemCom system is defined and obtained through adversarial convex optimization. Based on this semantic error, the parameters of the optimized robust encoder are updated from $\bm{\phi}$ to $\bm{\phi}^{\prime}$ to encode normal latent space for transmission. The specific GAN robust inversion method is detailed in Subsection III-B.
  2. 2.
    Domain Adaptation: When transmitting out-of-distribution data, the lightweight single-layer $g_{\bm{\omega}}$ and $d_{\bm{\nu}}$ are exploited for one-shot fast and adversarial domain adaptation learning. The learned parameters $\bm{\omega}$ will be seamlessly transmitted to the receiver along with the data for latent space transformation, and the decoder will ultimately output semantically consistent data. The specific implementation can be found in Subsection III-C.
  3. 3.
    Low-Latency Channel Denoising: Assuming that the CSIs are known, EECD is proposed to distill LDM from a multi-step denoising process into one setp, thereby reducing the computational complexity of online sampling during real-time communication. The detailed wireless channel modeling, training and sampling approaches of the latent channel denoising DM, and one-step real-time channel denoising algorithm are elucidated in Section IV.
These advancements open up a range of potential applications, including real-time video streaming, remote healthcare monitoring, and intelligent transportation systems, where low-latency and high-quality communication is crucial. Moreover, the ability to effectively manage semantic ambiguities and wireless channel noise further positions this system as a valuable solution in IoT environments and augmented reality applications, where accurate and timely semantic information exchange is essential.
III Deep JSCC for data with unknown errors and distributions